Ax, Ay, Dx, Dy, Md?
Determine the magnitude and direction of all support reactions for the pin-jointed structure shown below. 450 lb
Answer #1

Reacheus 450 lblft 41 & Fnzo; Ax B Ant Dazo < Fy=0) Ay 450 X 12 -+ Dy=0 41 V Ay+Ay = 5400 for member BC Dö 2 Mc=0; B Bx 3 ByŹ Fazo, Ax-Bn=0 An=Bx An= 5400 b () < Fyzo; Ay- 450x12-By zo by = 5400+ By = 5400-4050 My = 1350 lb (1) from equation i); Da=